OpenStudy (anonymous):

i not sure if i did this right. help please :) multiply using foil (10+square root 6)(10-square root 3) after using foil i got 100+17.32+24.49-4.24=137.57???

OpenStudy (anonymous):

help @hartnn

OpenStudy (anonymous):

someone just help :/

OpenStudy (anonymous):

It is not correct

OpenStudy (anonymous):

ok can you help me and explain how to get the correct answer

OpenStudy (anonymous):

\[ \left(10+\sqrt{x}\right) \left(10-\sqrt{y}\right)= 100 + 10 \sqrt x -10 \sqrt y - \sqrt {xy} \\ \] Replace x by 3 and y by 6 and use your calculator.

OpenStudy (anonymous):

ok so that is the equation we use every time we have a problem like this?

OpenStudy (anonymous):

Yes, in my previous post x=6 and y=3

OpenStudy (anonymous):

is the answer 82.43

OpenStudy (anonymous):

No

OpenStudy (anonymous):

i plunged in the numbers why is it not working?

OpenStudy (anonymous):

how about 88.58

OpenStudy (anonymous):

\[ \sqrt 6=2.44949\\ \sqrt 3= 1.73205\\ \sqrt {18}= 4.24264\\ \]

OpenStudy (anonymous):

ok well when i plugged in the x and ys this time i got 88.58

OpenStudy (anonymous):

100 + 10 (2.44949) -10 (1.73205) - 4.24264=102.932

OpenStudy (anonymous):

ok so are you suppose to get the square root of 6 3 and 18 before you plug them in?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank you for explaining that can you help me with one more question?

OpenStudy (anonymous):

Post is as a new problem and let me know when you do

OpenStudy (anonymous):

ok sounds good

OpenStudy (anonymous):

it is done on the top

OpenStudy (anonymous):

it is about the third from the top

OpenStudy (anonymous):

did you find it?